OpenStudy (anonymous):

ok simplify 2square root50 +13 square root 200 show work

OpenStudy (anonymous):

Find the factors that can have the square root taken of them in the numbers under the square root sign. 50 has 2,5,10,25,50,and 1 as multiples. Which ones can have the square root taken? Same with 200. That's the first step.

OpenStudy (anonymous):

\[2\sqrt{50}=2\sqrt{25\times 2}=2\times 5\sqrt{2}=10\sqrt{2}\] do the similar thing with 200

OpenStudy (anonymous):

Satellite has the better setup

OpenStudy (anonymous):

ok now how do i do the \[13\sqrt{200}\] part?

OpenStudy (anonymous):

Find the multiples of 200 that can have the square root taken of them. What are the multiples of 200?

OpenStudy (anonymous):

what brandon said. look for perfect squares inside 200. i think there is a pretty big obvious one

OpenStudy (anonymous):

144? im not sure what you mean

OpenStudy (anonymous):

2square root50 +13 square root 200 = 2sqrt(2*5*5) + 13* sqrt(2*2*2*5*5) Now if u have x*x (like 5*5) u can take it out from under the radical as x (5)

OpenStudy (anonymous):

The factors in 200 are numerous, but pick the ones that are perfect squares. The big one that satellite mentioned is 100. 100 goes into 200 two times and the square root can be taken out of 100. Multiply the number outside the square root (13) by the square root of 100 and leave the 2 inside the square root. Do you see how to do this?

OpenStudy (anonymous):

\[140\sqrt{2}\] this is the answer i got is it correct

OpenStudy (anonymous):

\[13\sqrt{200} \] \[13\sqrt{100*2}\] \[13*\sqrt{100} \sqrt{2}\] \[13*10\sqrt{2}\] See?

OpenStudy (anonymous):

You're close!

OpenStudy (anonymous):

because i had to add \[10\sqrt{2}+10*13\sqrt{2}\]

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

ok thanks :D